Diana wants to create a rectangular chocolate bar for their mini sweets store for her friend, Daisy, and for her treatment. She plans on making 500 chocolate bars that she wants to sell at $5 each. Given that the dimension of each chocolate bars are (6d+9) cm. and (4d+1) cm. in terms of length and width, respectively, and the bar's total perimeter is 278 cm. Find the value of d, the area of 1 chocolate bar, and the length of its diagonals(diagonals can be name randomly).
Answer by ikleyn(52781)

.

To find the value of d, write and solve this equation for the perimeter


    (6d+9) + (4d+1) + (6d+9) + (4d+1) = 278.


Simplify and solve for d


    20d + 20 = 278

    20d      = 278 - 20 

    20d      = 258

      d      = 258/20 = 12.9 cm


Having this value of d,  do the rest of the assignment on your own.


It is simple arithmetic.


To find the diagonal, use the Pythagorean theorem.

------------

Having your job partly done, boldly go forward from this point.